The horizontal and vertical components of a force are respectively 20 N and 30 N. Calculate the magnitude and direction of the force.
1
Expert's answer
2020-10-12T11:40:42-0400

The magnitude of the force can be found using the Pythagoras' Theorem.

Magnitude of the force "= \sqrt{ 20^{2} + 30^{2}} = \sqrt{1300} = 36.05N"

For the direction of the force, "\theta = tan^{-1}(\frac{vertical component}{horizontal component}) = tan^{-1}(\frac{30}{20}" ")"

"\theta = 56.31^{o}"

Therefore, the magnitude of the force is "36.05 N" and direction angle of "56.31^{o}"